SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

Minutes of the Trigger Operation Meeting (July 11, 2006)

Attendees: Rainer Bartoldus, Fang Fang, Selina Li, Rahmat Rahmat, Dan Walker

Operations

  • No new operational problems with the TRG.

EMT

  • Dan had trouble looking up the EMT deadtime from ambient database. Dan will learn how to obtain VMON data from the ODF people. Rainer mentioned that we should be able to predict the deadtime contribution from EMT for different suppression factor after the ROM0 bug was fixed. We will make different configuration alias for different suppression factor for ease of switching back and forth.

DCT/DCZ

  • The cooling radiator has been removed from the old DCT rack at IR2 in preparation for Henry Band's bubbler installation at rack 23. Ray Rodriguez is getting ready to remove the fiber splitter for TSF-Y, which will probably happen sometime next week during a ROD when we are not taking data. Some couplers (replacing the splitter) will be used to connect the fiber from DCH to the new TSF-Y interface board.

L3

  • We still need offline validation for the latest online release which hasn't been installed at IR2 yet. Rainer mentioned that Arik can probably help with this validation.

Data Quality and Monitoring

  • Rahmat observed about 5% difference in L3 lumi vs offline lumi. Rainer pointed out that this is expected. There is a correction to the L3 lumi to compare with the offline lumi. It was used to be ~6.8%, but now 7.4%.
  • Rahmat will have more lumi ratio strip charts to show for next week's meeting. Rainer pointed out that we would like to have the lumi strip charts in the official production release as soon as possible before the end of run 5b; otherwise we won't be able to see these strip chart till the beginning of next year when data-taking starts again.

Trigger Configuration

  • Selina showed the trigger statistics plots for some two-photon physics samples. Rainer will come up with more detailed instruction to show people how to use the Trg-statistics tool to run on their analyses.

This page is created by Selina Li (selina@slac.stanford.edu); however, you are welcome to edit and/or correct above items, specially for those who attended the meeting. Thanks.